It requires 0.86 aJ to ionize a single gas phase Li atom. How much energy is required to ionize 6.022 × 10^23 gas phase Li atoms?
Answer #1

Energy for 1 atom = 0.86 aJ

= 0.86*10^-18 J

So,

for 6.022*10^23 atoms,

energy = 6.022*10^23 * 0.86*10^-18 J

= 5.18*10^5 J

Answer: 5.18*10^5 J